if 1g of carbon contains xatoms,what will be the no. of atoms in 1g of magnesium
 

Grade:9

2 Answers

Darpan Verma
6 years ago
atomic mass of carbon =12 u.So, 12g of carbon = 12x atomsAlso, 12g of carbon = 6.022 ×10²³Therefore, 12x = 6.022 × 10²³Atomic mass of magnesium = 24 u.So, 24g of magnesium = 6.022 × 10²³We know that 6.022 × 10²³ = 12xNow, 24g of magnesium = 12xHence, 1g of magnesium = x/2 atomsYOUR ANSWER IS X/2 ATOMS.
praveen kumar
6 years ago
no of atoms depends upon the moles of the compound so,
moles of carbon=1/12 moles then number carbon atoms = (1/12)*avagadro number=x
moles of mg=1/24 then no of atoms mg =(1/24)*avagadro number=x/2
there fore the answer for your question is x/2